OpenStudy (jenniferjuice):

Solve for x: |x + 2| + 16 = 14

OpenStudy (jhannybean):

It's the same method for this one...instead of |x| we have | x+2 | so you know the out come is going to be 2 values. First isolate |x+2| to one side. |x+2| = 14-16 |x+2| = -2 Opening up the absolute value you'll have x+2 = -2 or x+2 = 2 Evaluate x for both of these functions.

OpenStudy (jenniferjuice):

x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (jenniferjuice):

i got zero and zero

OpenStudy (jenniferjuice):

?

OpenStudy (jenniferjuice):

@Jhannybean

OpenStudy (jenniferjuice):

okay thankyou

OpenStudy (jhannybean):

wait a minute, no....

OpenStudy (jenniferjuice):

wait whaaaaa

OpenStudy (jhannybean):

x+2= -2 x = -2 - 2 x+2 = 2 x = 2 - 2

OpenStudy (jenniferjuice):

now what?

OpenStudy (jhannybean):

Solve for x, i think you made an error. It's actually not 0, 0 ..

OpenStudy (jenniferjuice):

x+2 = -2 -2 -2 x= -4

OpenStudy (jenniferjuice):

yeah i messed up there

OpenStudy (jhannybean):

good.

OpenStudy (jhannybean):

so your values are?...

OpenStudy (jenniferjuice):

x = −4 and x = 0

OpenStudy (jhannybean):

yay! :D